OpenStudy (anonymous):

what is the mid point of a segment with the endpoints A(5,1) B(-3,7)

OpenStudy (anonymous):

Add the x coordinates of both the coordinate pair and divide it by 2.. Similarly do for y.. Can you do it??

OpenStudy (anonymous):

not really.

OpenStudy (anonymous):

is it (1,3)?..

OpenStudy (anonymous):

\[\huge (x, y) = (\frac{5-3}{2}, \frac{1+7}{2})\] Solve this..

OpenStudy (anonymous):

Very much close.. (1 , 4), I think..

OpenStudy (anonymous):

the answer only has the choice (0,4) and (1,4) @waterineyes

OpenStudy (anonymous):

So it is the later one.. (1. 4) Isn't it??

OpenStudy (anonymous):

I mean (2,4) not 1.4 @waterineyes

OpenStudy (anonymous):

No it is not 1.4 it is (1, 4) I think our answer is right.. Check if there are more choices..

OpenStudy (anonymous):

Can you attach the paqe where this question is?? I want to see it..

OpenStudy (anonymous):

those are the only two

OpenStudy (anonymous):

besides 1,3 and -1,-2

OpenStudy (anonymous):

Use the sroll bar may be you will get other two..

OpenStudy (anonymous):

heres a picture

OpenStudy (anonymous):

Show that one..

OpenStudy (anonymous):

These are high sized images... My net speed is very slow.. It will take time to be opened..

OpenStudy (anonymous):

Your eyes are good na??

OpenStudy (anonymous):

it is -1 and not 1.. \[\huge (x, y) = (\frac{5-3}{2}, \frac{7-1}{2})\] Now solve it..

OpenStudy (anonymous):

(2/2,6/2)

OpenStudy (anonymous):

Yes then you got the answer or not??